a)Calculate the potential for the reaction below at 25ºC under the following conditions:
[MnO?¯ ]=0.0200M
[Ni+2 ]=2.50M
pH = 2.000

2 MnO?¯ (aq) + 8 H+ (aq) + 3 Ni (s) ? 2 MnO? (s) + 4 H?O (l) + 3 Ni+2 (aq)
Answer #1

Eocell = Eocathode - Eo anode

          = 1.68 + 0.26

         = 1.94 V

Q = [Ni+2]^3 / [MnO4-]^2 [H+]^8

Q = (2.50)^3 / (0.02)^2 (10^-2)^8

Q = 3.91 x 10^20

Ecell = Eo cell - 0.05916 / n log Q

          = 1.94 - 0.05916 / 6 * log (3.91 x 10^20)

          = 1.74 V

Ecell = 1.74 V